An industry consists of three firms with sales of $225,000, $485,000, and $320,000.

a. Calculate the Herfindahl-Hirschman index (HHI). Instruction: Enter your response rounded to the nearest integer.

b. Calculate the four-firm concentration ratio (C4).

c. Based on the FTC and DOJ Horizontal Merger Guidelines described in the text, is the Department of Justice likely to attempt to block a horizontal merger between two firms with sales of $225,000 and $320,000?
Answer #1

Answer:-

(A). Calculate the Herfindahl - Hirschman index (HHI):-

Total sales = 225,000 + 485,000 + 320,000 = 10,30,000

Herfindahl-Hirschman index (HHI)

= (225,000 / 10,30,000 *100)^2 +(485,000 / 10,30,000 * 100)^2 + (3,20,000 / 10,30,000 * 100)^2

= 477.19 + 2217.22 + 965.23

= 3659.64

(B). Calculate the four-firm concentration ratio (C4)

= (225,000 / 10,30,000) + (485,000 / 10,30,000) + (320,000 / 10,30,000)

= 0.2184 + 0.4709 + 0.3107

=100.00%

(C). If the two firms with sale 225,000 and 320           ,000 merge then the HHI will be

HHI = (225,000 + 320,000 / 10,30,000 * 100)^2 + (485,000 / 10,30,000 * 100)^2

    = (5,45,000 / 10,30,000 * 100)^2 + (485,000 / 10,30,000 * 100)^2

    = 2799.76 + 2217.22

    = 5016.98

Thus, the HHI after merger increase to 5016.98. Therefore, HHI after merger increase by 1357.34